OpenStudy (experimentx):

Let \( f(x) \) and \( g(x) \) be two increasing and continuous functions on interval \( (a, b) \). Check weather the following integral is positive or not. \[ \int_a^b f(x) g(x) dx \] Also let \[ \int_a^b f(x) g(x) dx =0 \] define the scalar product of two functions \( f(x) \) and \( g(x) \)

OpenStudy (experimentx):

@hidy give it a try.

OpenStudy (anonymous):

okay

OpenStudy (kc_kennylau):

NEI; not enough information

OpenStudy (kc_kennylau):

f(x)g(x) can be below the x-axis or above the x-axis

OpenStudy (kc_kennylau):

f(x) and g(x) can be negative or positive

OpenStudy (anonymous):

yes, but in the giving integral you choosed to work in one dimension IR what if the integral is defined like this way :\[\int\limits\int\limits f(x)g(y)dxdy\]

OpenStudy (anonymous):

remember that we want to proove it for the bilinear form B(x,y)

OpenStudy (kc_kennylau):

Sacred faeces this is getting intense

OpenStudy (experimentx):

looks like I didn't not put up enough information!! ... this was to check the claim by @hidy "if we compose a bilinear form with two increasing functions do we get a positive biliear form"

OpenStudy (anonymous):

The integral could be positive or negative

OpenStudy (anonymous):

f(x) = x^2 and g(x) = x - 1, on [0,1], the integral is negative or f(x) = x^2 and g(x) = x , on [0,1] the integral is positive

OpenStudy (anonymous):

the biliner form dosen't necessary define a scalar product

OpenStudy (experimentx):

I don't think we need to define the scalar product that way ... @hidy ... probably you would do that for smooth functions of two variables i.e. f(x,y), g(x,y) on some interval (a,b)x(c,d). just stick with single integral for one variable. If there is counterexample then, your claim is invalid.

OpenStudy (experimentx):

probably I should make the interval (a,b) such that \[\int_a^b f(x) g(x) dx = 0 \] if f(x) and f(x) are linearly independent.

OpenStudy (experimentx):

well ... you are definitely right, but we need a map there that maps two functions into field. couldn't think any better than scalar product.

OpenStudy (anonymous):

and what can that give us ??

OpenStudy (anonymous):

yes I see

OpenStudy (experimentx):

i kinda feel hard to define linear independence ... to make sourwing claim invalid.

OpenStudy (anonymous):

I got an idea but I don't know how to use it we know that we can associate a quadratic form to a bilnear form that satisfies B(x,x)=q(x) and q is always positive

OpenStudy (experimentx):

still you need a map \[ B:L^2\times L^2 \to \Bbb R \]

OpenStudy (anonymous):

yeah :(

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Quadratic_form the first example you see is for vector space (x,y) ... this x, and y represents numbers.

OpenStudy (anonymous):

check this out The associated bilinear form of a quadratic form q is defined by \[ b_q(x,y)=\tfrac{1}{2}(q(x+y)-q(x)-q(y)) = x^\mathrm{T}Ay = y^\mathrm{T}Ax. \]

OpenStudy (anonymous):

so obviously the bilinear form is not always positive we should have the condition q(x+y)>q(x)+q(y)

OpenStudy (experimentx):

no it won't ...

OpenStudy (anonymous):

so I guess that's it we cant' go farther but if I find something I'll show you thanks for your help :)

OpenStudy (experimentx):

yw ... i hope someone answers it, algebra isn't my cup of tea.

OpenStudy (experimentx):

yw ... i hope someone answers it, algebra ain't my cup of tea.

OpenStudy (anonymous):

no no it was an amazing effort I'm impressed :)